OpenStudy (anonymous):

is 1 over 11 the slope of 8,8 and 7,-3

OpenStudy (gorv):

1 over 11= 1/11

OpenStudy (gorv):

slope =(8-(-3))/(8-7)=11/1

OpenStudy (gorv):

welcome to OS :)... u r new here so any Q related to os ask here http://openstudy.com/study#/groups/OpenStudy%20Ambassadors